# DEIO: Deep Event Inertial Odometry

### v2

@@ II-A Traditional Approaches for Event-based SLAM
The first purely event-based VO was presented in [11], which performed real-time event-based SLAM through three decoupled probabilistic filters that jointly estimate the 6-DoF camera pose, 3D map of the scene, and image intensity. However, it is computationally expensive and requires GPU to achieve real-time performance. EVO [1] presents a geometric 2D-3D model alignment, utilizing the 2D image-like event representation and the 3D map reconstructed from EMVS[12]. It can perform on small-scale scenes, but it is highly parameter-sensitive and requires to run in the scene that is planar to the sensor, up to several seconds, for bootstrapping the system. ESVO [2] is the first stereo event-based VO method to estimate the ego motion through the 2D-3D edge registration on time surface. However, it barely operates in real-time in low resolution event camera and also faces limitations due to rigorous initialization as well as unreliable pose tracking. What’s more, some approaches reply on a pre-build photometric 3D map [13] or restrict the motion type to rotation-only [14]. However, most of them are prone to failure in challenging scenarios (such as fast motion), which negates the unique advantages of event cameras.
To enhance the robustness of purely event-based VO, existing event-based SLAM methods have demonstrated good performance by incorporating additional sensors. Notably, event-inertial integration is a widely used approach to address the limitations of event-only SLAM, which provides scale awareness and continuity of estimation with minimal setup requirements. A monocular event-inertial odometry (EIO) is presented in [10], which employs the event-corner features with IMU measurement to deliver real-time and accurate 6-DoF state estimation. What’s more, this EIO framework can provide good global consistency through loop closure of the event-based representation. Ultimate-SLAM [3] and PL-EVIO [4] investigate the complementary nature of events and images to present an event-image-IMU odometry (EVIO). Chamorro et al. [15] integrate the event-based line feature with the IMU to achieve high ultra-fast pose estimation. ESVIO [16] proposed the first stereo EIO and EVIO framework to estimate states through temporally and spatially event-corner feature association. EVI-SAM [17] introduced the first event-based hybrid pose tracking framework, merging the feature-based and direct-based methods. Additionally, it consistently delivers excellent mapping results in a variety of challenging scenarios, demonstrating performance comparable to learning-based and NeRF-based mapping methods. Despite such remarkable success, current event-based SLAM systems still fall short of the robustness required for certain real-world applications, and their performance has hit a bottleneck, this directs toward the exploration of deep learning approaches.
@@ II-B Learning-based Approaches for Event-based SLAM
DH-PTAM [22] proposed a stereo event-image parallel tracking and mapping system that replacing hand-crafted features with learned features. Klenk, et al. [9] extends the DPVO [23] to accommodate the event modality through voxel-based representation, which are able to demonstrate generalization from simulation to several real-world event-based benchmarks. It shows good performance compared to both classical and deep-learning baselines. However, their results reveal a significant scale loss/ambiguity, and they requires the careful selection of the input data. Pellerito, et al. [24] introduces an end-to-end VO which also builds upon DPVO using complicated feature encoders to fuse asynchronous events with image data. However, it requires a minimum of 1.2 million events between each pair of frames and discards additional frames if this requirement is not met. More importantly, it failed to evaluate their methods under fast motion or HDR scenes, and could only operate in very limited amount of real event data without challenging situations.
Our framework is closely related to [23, 9] but extends these works by integrating trainable event representations with IMU through our learning-optimization-combined strategy to improve the performance and robustness. Additionally, it is important to note that a visual-only (event-only) system, such as DPVO [23] and DEVO [9] has inherent limitations, failures can manifest in numerous ways, such as in low-texture environment, divergence in the optimization algorithm, scale ambiguity, and etc. Nevertheless, there is a absence of an implementation that incorporates learning-based event SLAM with IMU fusion framework.

### v3

@@ II-A Traditional Approaches for Event-based SLAM
The first purely event-based VO was presented in [11], which performed real-time event-based SLAM through three decoupled probabilistic filters that jointly estimate the 6-DoF camera pose, 3D map of the scene, and image intensity. However, it is computationally expensive and requires GPU to achieve real-time performance. EVO [1] presents a geometric 2D-3D model alignment, utilizing the 2D image-like event representation and the 3D map reconstructed from EMVS[12]. ESVO [2] is the first stereo event-based VO method to estimate the ego motion through the 2D-3D edge registration on time surface. What’s more, some approaches reply on a pre-build photometric 3D map [13] or restrict the motion type to rotation-only [14]. However, most of them are prone to failure in challenging scenarios (such as fast motion).
To enhance the robustness of purely event-based VO, existing event-based SLAM methods have demonstrated good performance by incorporating additional sensors. Notably, event-inertial integration is a widely used approach to address the limitations of event-only SLAM, which provides scale awareness and continuity of estimation with minimal setup requirements. Zhu et al. [15] proposed the first event-inertial odometry (EIO) which fuses events with IMU through the Extended Kalman Filter. A monocular EIO is presented in [10], which employs the event-corner features with IMU measurement to deliver real-time and accurate 6-DoF state estimation. Ultimate-SLAM [3] and PL-EVIO [4] investigate the complementary nature of events and images to present an event-image-IMU odometry (EVIO). Chamorro et al. [16] integrate the event-based line feature with the IMU to achieve high ultra-fast pose estimation. ESVIO [17] proposed the first stereo EIO and EVIO framework to estimate states through temporally and spatially event-corner feature association. EVI-SAM [18] introduced the first event-based hybrid pose tracking framework, merging the feature-based and direct-based methods. Additionally, it consistently delivers excellent mapping results in a variety of challenging scenarios, demonstrating performance comparable to learning-based and NeRF-based dense mapping methods. Despite such remarkable progress, current event-based SLAM systems still fall short of the robustness required for certain real-world applications, and their performance has hit a bottleneck, this directs toward the exploration of deep learning approaches.
@@ II-B Learning-based Approaches for Event-based SLAM
DH-PTAM [23] proposed a stereo event-image parallel tracking and mapping system that replacing hand-crafted features with learned features. Klenk, et al. [9] extends the DPVO [24] to accommodate the event modality through voxel-based representation, which are able to demonstrate great generalization from simulation to several real-world event-based benchmarks. It shows good performance compared to both classical and deep-learning baselines. RAMP-VO [25] introduces an end-to-end VO which also builds upon DPVO using feature encoders to fuse events and image data. It achieves robust zero-shot transfer performance to real data despite being trained only on synthetic dataset.
Our framework is closely related to [24, 9] but extends these works by integrating trainable event-based optical flow with IMU through our learning-optimization-combined strategy to improve the performance and robustness. It is important to note that visual-only (or event-only) systems have inherent limitations, with failures manifesting in various ways, such as in low-texture environments, optimization algorithm divergence, and scale ambiguity. Visual-IMU integration is the most common solution to address these limitations, which providing scale awareness and continuous estimation with minimal setup. Nevertheless, there is a absence of implementation that combines learning-based event SLAM with IMU.